# From #Dr00gtiktok to #harmreduction: Exploring Substance Use Hashtags on TikTok

### v1

@@
Abstract
The rise of TikTok as a primary source of information for youth, combined with its unique short-form video format, creates urgent questions about how substance use content manifests and spreads on the platform. This paper provides the first in-depth exploration of substance use-related content on TikTok, covering all major substance categories as classified by the Drug Enforcement Agency. Through social network analysis and qualitative coding, we examined more than 2,333 hashtags across 39,509 videos, identified 16 distinct hashtag communities and analyzed their interconnections and thematic content. Our analysis revealed a highly interconnected small-world network where recovery-focused hashtags like “#addiction”, “#recovery”, and “#sober” serve as central bridges between communities. Through manual coding of 351 representative videos, we found that Recovery Advocacy content (33.9%) and Satirical content (28.2%) dominate, while direct substance depiction appears in only 26% of videos, with active use shown in just 6.5% of them. This suggests TikTok functions primarily as a recovery support platform rather than a space promoting substance use. We found strong alignment between hashtag communities and video content, indicating organic community formation rather than attempts to evade content moderation. Our findings inform how platforms can balance content moderation with preserving valuable recovery support communities, while also providing insights for the design of social media-based recovery interventions.
Introduction
Social media platforms have become the primary lens through which young people view and understand the world, with TikTok emerging as a dominant force among youth - reaching 63% of teens, with 17% reporting near-constant use (Vogels, Gelles-Watnick, and Anderson 2023). More concerning is that TikTok has become a primary news source for this demographic, often replacing Google for information-seeking (Walker and Matsa 2023).
Figure 1: Examples of TikTok videos that can be found in the substance use communities based on hashtag use
This shift in information consumption patterns takes on particular urgency when considering content related to illicit substances. While overall drug use among 8th, 10th, and 12th graders showed promising declines during COVID-19 social distancing (National Institute on Drug Abuse 2022), this drop in usage may have masked how drug-related discussions were shifting online. Social media platforms can rapidly disseminate trends and behaviors through psychological processes like social comparison and fear of missing out (FOMO) (Montag, Yang, and Elhai 2021), with adolescents often imitating influencers and peers they follow. Although most social media platforms use content moderation, their rapid and viral nature can make it challenging to effectively regulate exposure to harmful content (Boyd 2014; Lookingbill and Le 2024), potentially amplifying dangerous trends. Therefore, understanding such emerging trends on social media platforms such as TikTok, Instagram, and Reddit (Rutherford et al. 2022a, 2023; Tan and Weinreich 2021; Bouzoubaa, Young, and Rezapour 2024) becomes more critical. For instance, the “xandemic” between 2016-2018, characterized by high school students documenting the recreational use of Xanax to the point of blackout, exemplifies how social media trends can blur the lines between casual and harmful behaviors and even normalize dangerous substance use behaviors (Fuller et al. 2024). Similarly, platform-specific hashtag “#pingtok” - a portmanteau of the slang term for an MDMA pill, pinger, and TikTok - which has accumulated over 26,000 posts at the time of writing, demonstrates how substance-related content can become embedded in platform culture, with its own specific terminology and community. These trends are especially concerning given robust evidence that early exposure to substance use significantly increases the risk of developing substance use disorders (SUD) in adulthood (McCabe et al. 2022; Volkow and Wargo 2022; Odgers et al. 2008).
@@ Introduction
Previous research on substance-related content on TikTok has typically focused on either specific substances such as cannabis or nicotine (Rutherford et al. 2023, 2022a; Morales, Fahrion, and Watkins 2022; Tan and Weinreich 2021) or isolated trending phenomenon like “#pingtok” (Whelan, Noller, and Ward 2024). While some researchers have examined hashtag networks and content moderation strategies around other sensitive topics like eating disorders (Bickham et al. 2024) and how hashtags create communities around shared experiences (Eriksson Krutrök 2021), there remains limited understanding of how hashtags shape substance-related discourse on TikTok. Moreover, no studies have comprehensively examined how substance-related content manifests across TikTok’s ecosystem of communities, particularly through its short-form video format and hashtag-driven discovery system (Abidin 2020). To address this gap, we draw from the Drug Enforcement Agency’s comprehensive drug classification system and examine content from every major substance category using a mixed-methods investigation of TikTok content. Our research questions are:
@@
  • RQ3: What are the effects of substance visibility on topic distribution and engagement metrics?
For this purpose, we developed a list of 585 seed hashtags related to substance use and collected more than a thousand videos for each of them, resulting in a dataset of 504,294 hashtags and 272,910 videos  [^1] Datasets, including seed hashtags, and code are available at https://github.com/social-nlp-lab/explore˙tiktok˙su. Videos are not directly provided but can be accessed via URLs.. After filtering relevant hashtags, we identified 2,333 exemplary hashtags covering the broad categories of substance use, and 351 videos for qualitative coding. Using a combination of network analysis of hashtag co-occurrence and in-depth content analysis, our analysis showed a highly interconnected small-world network where recovery-focused hashtags like “#addiction”, “#recovery”, and “#sober” serve as central bridges between communities. Within these communities, we found that Recovery Advocacy content (33.9%) and Satirical content (28.2%) dominate, while direct substance depiction appeared in only 26% of videos, with active use shown in just 6.5%. This suggests TikTok functions primarily as a recovery support platform rather than a space promoting substance use.
Through this work, we contribute to social computing and substance use literature by providing the first comprehensive mapping of substance-related content on TikTok. Our findings can inform platform policies around content moderation strategies that preserve supportive recovery communities while better identifying potentially harmful content, guide the development of social media-based recovery support interventions, and help design youth substance use prevention efforts that account for how substance-related content actually manifests on TikTok.
Related Works
@@
As digital platforms become ubiquitous, understanding their impact, especially on young people’s health behaviors regarding substance use, is more crucial than ever. On the one hand, social media emerged as a medium for facilitating sensitive discussions and disclosure of lived experiences, including mental health and eating disorders (Rutherford et al. 2022a; Bickham et al. 2024; Bouzoubaa et al. 2024). On the other hand, several studies have examined the relationship between online peer groups, social media homophily, and substance use, finding positive correlations between substance use and social media engagement. For example, Willoughby et al. (2024) found that young adults reported an increased intention to use cannabis after exposure to pro-cannabis messages online. Similarly, Sun et al. (2023) showed that exposure to e-cigarette advertisements on social media platforms is associated with adolescents’ e-cigarette usage in later years. Miller et al. (2021) examined how online and offline peer interactions influence substance use behaviors and found that online belonging and social media homophily predicted regular stimulant and opioid use among substance users. Hanson et al. (2013) have studied online discussions related to stimulant drug use on Twitter and found a high correlation between the popularity of such topics and final exam periods. Geusens et al. (2023) found that higher levels of alcohol and marijuana consumption predicted later social media posts about those substances, but posting did not predict later substance use. Sun et al. (2023) found a positive correlation between early exposure to e-cigarette advertisement and subsequent lifetime usage.
@@
Hashtags, a medium for interactions on platforms like Facebook or Twitter, have been studied comprehensively in the past (Mousavi and Ouyang 2021; Ince, Rojas, and Davis 2017; Small 2011; Rezapour et al. 2017). People can contribute to the narrative by hash-tagging their own personal experiences and thoughts (Yang 2016). At their core, hashtags facilitate content discovery and community creation by allowing users to contribute to the discussion through keyword tagging (Small 2011). Previous studies have examined substance use discourse across various social media platforms, contributing to our understanding of how substance use behaviors manifest online. For instance, on Instagram, investigating “#studydrugs” demonstrated a predominantly positive portrayal of stimulant use for academic performance enhancement (Petersen et al. 2021).
@@
On TikTok, content creators deliberately employ hashtags like “#fyp” (For You Page) to increase visibility and maximize audience reach within TikTok’s recommendation system (Abidin 2020). The platform’s ranking algorithms connect users based on hashtag engagement patterns, creating what Eriksson Krutrök (2021) calls “algorithmic closeness” - subcommunities united through shared content interests. This algorithmic dynamic shapes how substance-related discourse spreads and evolves on TikTok. For example, research examining TikTok hashtags like “#puffbar” has revealed concerning patterns of youth-oriented substance marketing and normalization (Tan and Weinreich 2021), while analysis of “#stonertok” and “#420vibes” showed the majority of cannabis-related videos positively portray substances (Rutherford et al. 2022b).
Our study advances existing research by employing a network-based and qualitative approach to explore the interactions and communities associated with substance-related hashtags on TikTok. While previous research has thoroughly documented individual communities, the dynamics of how these groups interact across the platform’s broader ecosystem remain understudied. Our work can help social media platforms like TikTok refine their policies to better distinguish between harmful content requiring removal and harm reduction information that should be preserved, contributing to more effective content moderation.
Methods
Data
@@ Methods
Seed Hashtag Selection: To develop an initial set of seed hashtags related to substance use for the data collection, we first identified hashtags previously used in studies of substance-related content on social media platforms (Petersen et al. 2021; Russell et al. 2022; Rutherford et al. 2023, 2022b; Singh and Wu 2021) as well as searching for hashtags on the platform associated with substance names (both formal and slang) listed on the Drug Enforcement Agency’s Fact Sheet (United States Drug Enforcement Administration 2025). Using this set, we conducted iterative searches on TikTok to identify additional relevant hashtags by examining co-occurring hashtags within the content. This process continued until we reached saturation, resulting in a final set of 585 substance-use-related hashtags. [^2] The list will be shared upon the paper’s acceptance.
@@
Data Collection: Using the Unofficial TikTok API [^3] https://github.com/davidteather/TikTok-Api, we collected data for around 1,000 algorithmically served videos for each of the 585 seed hashtags. This resulted in a dataset of 275,291 videos and a total of 504,294 unique hashtags.
Figure 2: Data selection and pre-processing pipeline to create the filtered set.
@@
Data Preprocessing and Network Refinement: To enhance the feasibility of our analysis and reduce noise from potentially irrelevant content, we filtered the data to include only videos that featured at least two of the seed hashtags. This filtering step resulted in a refined dataset of 39,602 videos and 99,727 unique hashtags (Figure 2). To further clean the data, we constructed a co-occurrence network where nodes represent individual hashtags and edges represent their co-occurrence in videos. More specifically, edge weights were defined by the frequency with which two hashtags appeared together in videos. We then employed a two-stage filtering process to focus on the most relevant and interconnected hashtags. First, we removed hashtags with low connectivity by filtering out nodes in the bottom 25th percentile of degree centrality. This step eliminated peripheral hashtags with minimal connections to the broader substance use discourse, removing those that were incidentally rather than substantively related to substance use discussions. Next, we applied k-core decomposition ($k=10$) to identify the densely connected core of the network (Seidman and Foster 1978). K-core decomposition recursively removes nodes with degrees less than $k$ until all remaining nodes have at least $k$ connections, ensuring a cohesive network structure. This threshold was chosen based on exploratory analysis, which revealed that $k=10$ produced the most structurally cohesive subset of hashtags while retaining a sufficient number of nodes and edges to analyze the discourse. The resulting refined network of hashtags contained 2,333 nodes and 46,990 edges, representing the most interconnected hashtags in substance use-related content.
@@
Table 1: Categories for semantically classifying hashtags from TikTok content, their definitions, and example hashtags.
Substance Use Hashtags Network
@@
Community Detection: We first applied established community detection algorithms including Clauset-Newman-Moore (Clauset, Newman, and Moore 2004) and Louvain method (Blondel et al. 2008), but these approaches did not yield clusters that meaningfully captured substance-related discussions. Therefore, we used a hybrid approach combining manual and automated methods to identify thematic groups within the hashtags. We first manually annotated a sample of 251 hashtags and iteratively established thematic categories. As a result, we identified 16 categories representing topics related to substance use and consumption, identity, and community, health and emotion as well as awareness and advocacy. In addition, we created a Misc category for general hashtags, not related to substance use content. The final set of categories and their definitions are shown in Table 1. These categories were then used to prompt GPT-4o (OpenAI 2024) to classify hashtags. When evaluated against human-labeled data, our best-performing prompt achieved an accuracy of $89.36\%$, demonstrating strong reliability in the thematic categorization. We then used this prompt to classify the full set of 2,333 hashtags. The resulting labeled hashtags were mapped back onto the network, creating themed communities that reflect different aspects of substance use discourse on TikTok.

### v2

@@
Abstract.
TikTok has emerged as a major source of information and social interaction for youth, raising urgent questions about how substance use discourse manifests and circulates on the platform. This paper presents the first comprehensive analysis of publicly visible, algorithmically surfaced substance-related content on TikTok, drawing on hashtags spanning all major substance categories. Using a mixed-methods approach that combines social network analysis with qualitative content coding, we examined 2,333 substance-related hashtags, identifying 16 distinct hashtag communities and characterizing their structural and thematic relationships. Our network analysis reveals a highly interconnected small-world structure in which recovery-focused hashtags such as #addiction, #recovery, and #sober serve as central bridges between communities. Qualitative analysis of 351 representative videos shows that Recovery Advocacy content (33.9%) and Satirical content (28.2%) dominate, while direct substance depiction appears in only 26% of videos, with active use shown in just 6.5% of them. These findings suggest that the algorithmically surfaced layer of substance-related discourse on TikTok is predominantly oriented toward recovery, support, and coping rather than explicit promotion of substance use. We further show that hashtag communities and video content are closely aligned, indicating that substance-related discourse on TikTok is shaped through organic community formation within platform affordances rather than widespread adversarial evasion of moderation. This work contributes to social computing research by showing how algorithmic visibility on TikTok shapes the organization of substance-related discourse and the formation of recovery and support communities.
TikTok, Substance Use, Social Media, Hashtags, Generative AI, Socio-technical System, Governance
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/25/06
1. Introduction
Social media platforms have become critical and contested spaces for health discourse, particularly around stigmatized topics like substance use (russell_using_2021; Johnson et al., 2025). TikTok’s algorithmic recommendation system has demonstrated serious harms: amplifying health misinformation (Yeung et al., 2022), promoting eating disorders (Blackburn and Hogg, 2024), and contributing to mental health crises among young users (Conte et al., 2025). Yet these same algorithmic affordances enable the formation of peer support communities around substance use recovery, harm reduction, and lived experience sharing (russell_using_2021; Goyal et al., 2025). This tension reflects broader dynamics of platform governance, in which algorithmic visibility simultaneously mediates risk, care, and community formation in stigmatized health contexts, yet remains underexplored for substance use discourse, where both risks and potential benefits are acute.
Figure 1. Examples of TikTok videos illustrating distinct substance-related content identified through hashtags.
Understanding how substance use content manifests on platforms like TikTok is critical for multiple stakeholders. Early exposure to substance use significantly increases the risk of developing substance use disorders (SUD) (McCabe et al., 2022; Volkow and Wargo, 2022; Odgers et al., 2008), making youth exposure a public health concern. However, platforms also serve people already living with SUD who may lack offline support systems (russell_using_2021; Bouzoubaa et al., 2024b; Goyal et al., 2025). Previous trends like the “xandemic” (2016-2018), characterized by high school students documenting the recreational use of Xanax to the point of blackout, and #pingtok, a portmanteau of the slang term for an MDMA pill, pinger, and TikTok, demonstrate how substance-related content can become embedded in platform culture (Fuller et al., 2024; rutherford_changes_2022). What remains unclear is how substance normalization and recovery-support dynamics coexist, intersect, and vary across substances and communities.
@@ 1. Introduction
Previous research on substance-related content on TikTok has typically focused on either specific substances such as cannabis or nicotine (rutherford_turnttrending_2023; rutherford_changes_2022; morales_nicotineaddictioncheck_2022; tanPuffBarHowTop2021) or isolated trending phenomena such as #pingtok (Whelan et al., 2024). While some researchers have examined hashtag networks and content moderation strategies around other sensitive topics like eating disorders (Bickham2024HiddenInPlainSight) and how hashtags create communities around shared experiences (eriksson_krutrok_algorithmic_2021), there is limited comprehensive knowledge of how substance-related discourse is organized across TikTok’s hashtag-driven ecosystem. In contrast to prior work that examines individual substances, isolated trends, or moderation outcomes, we take a platform-wide, cross-substance perspective that captures how substance discourse is structurally organized and socially enacted across TikTok’s hashtag ecosystem. To address this gap, we draw on an established drug classification system (United States Drug Enforcement Administration, 2025) in order to examine content from every major substance category on TikTok. Our research questions are:
@@
  • RQ3: How does substance visibility, operationalized through engagement and algorithmic surfacing, affect topic distribution and engagement metrics?
We address these questions through a mixed-methods analysis combining social network analysis of hashtag co-occurrence patterns with qualitative content coding. Analyzing 2,333 exemplar hashtags and coding 351 representative videos, we examine both the structural organization of substance-related discourse and the actual content users create. Our network analysis reveals a highly interconnected discourse where recovery-focused hashtags like #addiction, #recovery, and #harmreduction serve as bridges between thematic communities. Qualitative coding shows that Recovery Advocacy (33.9%) and Humor-based coping (28.2%) make up the majority of videos, while direct substance depiction appears in only 26% of videos, with active use shown in just 6.5%. Our findings suggest that substance-related content on TikTok, within the publicly visible, algorithmically surfaced layer accessible to observation, is predominantly oriented toward support rather than promotion or glamorization.
Through this work, we contribute to social computing and substance use literature by providing the first comprehensive mapping of substance-related content on TikTok across all major drug categories, revealing 16 distinct hashtag communities within a highly interconnected small-world network. Second, we demonstrate that hashtag communities exhibit organic thematic coherence rather than adversarial evasion of content moderation, suggesting these support communities form within rather than against platform affordances. Third, we show that recovery advocacy (33.9%), humor-based coping (28.2%), and educational information (13.1%) collectively make up most of the content, with active substance use depiction making up only 6.5% in algorithmically surfaced public discourse. Together, these contributions advance social computing research on algorithmic visibility, health discourse, and community formation under platform governance constraints. Finally, we derive design implications for content moderation strategies that must balance harm prevention with preservation of beneficial peer support communities. These findings inform platform governance approaches, the development of social media-based recovery interventions, and youth substance use prevention efforts that account for how substance-related content actually manifests on TikTok.
2. Related Work
@@
As digital platforms become ubiquitous, understanding their impact, especially on young people’s health behaviors regarding substance use, is more crucial than ever. On the one hand, social media emerged as a medium for facilitating sensitive discussions and disclosure of lived experiences, including mental health and eating disorders (rutherford_changes_2022; Bickham2024HiddenInPlainSight; Bouzoubaa et al., 2024a). On the other hand, several studies have examined the relationship between online peer groups, social media homophily, and substance use, finding positive correlations between substance use and social media engagement. For example, willoughby_exposure_2024 found that young adults reported an increased intention to use cannabis after exposure to pro-cannabis messages online. Similarly, sun_longitudinal_2023 showed that exposure to e-cigarette advertisements on social media platforms is associated with adolescents’ e-cigarette usage in later years. miller_online_2021 examined how online and offline peer interactions influence substance use behaviors and found that online belonging and social media homophily predicted regular stimulant and opioid use among substance users. hanson_tweaking_2013 have studied online discussions related to stimulant drug use on Twitter and found a high correlation between the popularity of such topics and final exam periods. geusens_first_2023 found that higher levels of alcohol and marijuana consumption predicted later social media posts about those substances, but posting did not predict later substance use. sun_longitudinal_2023 found a positive correlation between early exposure to e-cigarette advertisement and subsequent lifetime usage.
@@
Hashtags, a medium for interactions on platforms like Facebook or Twitter, have been studied comprehensively in the past (mousavi_detecting_2021; ince_social_2017; small_what_2011; Rezapour et al., 2017). At their core, hashtags facilitate content discovery and community formation by enabling users to collectively construct and navigate shared narratives through keyword tagging (Yang, 2016; small_what_2011). Previous studies have examined substance use discourse across various social media platforms, contributing to our understanding of how substance use behaviors manifest online. For instance, on Instagram, investigating #studydrugs demonstrated a predominantly positive portrayal of stimulant use for academic performance enhancement (petersen_studydrugspersuasive_2021).
@@
On TikTok, content creators deliberately employ hashtags like #fyp (For You Page) to increase visibility and maximize audience reach within TikTok’s recommendation system (abidin_mapping_2020). The platform’s ranking algorithms connect users based on hashtag engagement patterns, creating what eriksson_krutrok_algorithmic_2021 calls “algorithmic closeness” - subcommunities united through shared content interests. This algorithmic dynamic shapes how substance-related discourse spreads and evolves on TikTok. For example, research examining TikTok hashtags like #puffbar has revealed concerning patterns of youth-oriented substance marketing and normalization (tanPuffBarHowTop2021), while analysis of #stonertok and #420vibes showed the majority of cannabis-related videos positively portray substances (rutherford_getting_2022).
Our study advances existing research by employing a network-based and qualitative approach to explore the interactions and communities associated with substance-related hashtags on TikTok. While prior research has examined individual substances or discrete hashtag communities, the structural relationships and interactions among these communities within TikTok’s broader ecosystem remain underexplored. Rather than focusing on message content or sentiment alone, we examine how substance-related discourse is structurally organized and interconnected at the platform level. Our work provides empirical grounding for understanding algorithmic visibility, community formation, and the coexistence of harm and support within substance-related discourse, informing not only content moderation but also platform governance, intervention design, and the development of evidence-based recommendations for public health and recovery-oriented stakeholders.
3. Methods
Figure 2. Data selection and pre-processing pipeline to create the filtered set.
3.1. Data
3.1.1. Seed Hashtag Selection:
@@
To develop an initial set of seed hashtags related to substance use for the data collection, we first identified hashtags previously used in studies of substance-related content on social media platforms (petersen_studydrugspersuasive_2021; russell_analyzing_2022; rutherford_turnttrending_2023; rutherford_getting_2022; singh_sentiment_2021) as well as hashtags associated with substance names (including both formal and slang terms) listed on the Drug Enforcement Agency’s Fact Sheet (United States Drug Enforcement Administration, 2025). Using this set, we conducted iterative searches on TikTok to identify additional relevant hashtags by examining co-occurring hashtags within the content. This process continued until we reached saturation (no substantively new substance-related hashtags yielded), resulting in a final set of 585 substance-use-related hashtags. [^1] The list will be shared upon the paper’s acceptance.
3.1.2. Video Collection:
@@
Using the Unofficial TikTok API [^2] https://github.com/davidteather/TikTok-Api, we collected up to 1,000 publicly available, algorithmically surfaced videos for each of the 585 seed hashtags during the data collection period. This resulted in an initial, raw dataset of 272,910 videos. These videos contained a total of 504,294 unique hashtags (including both substance-related terms and general hashtags that co-occurred with them).
3.1.3. Network Construction and Refinement:
@@
To enhance the feasibility of our analysis and reduce noise from potentially irrelevant content, we filtered the data to include only videos that featured at least two of the seed hashtags, ensuring that retained videos reflected explicit engagement with substance-related discourse rather than incidental hashtag use. This filtering step resulted in a refined dataset of 99,727 unique hashtags (Figure 2). To further clean the data, we constructed a hashtag co-occurrence network where nodes represent individual hashtags and edges represent their co-occurrence in the same video. Edge weights were defined as the frequency with which two hashtags appeared together in videos.
We then employed a two-stage network refinement process to focus on the most relevant and interconnected hashtags. First, we removed hashtags with low connectivity by filtering out nodes in the bottom 25th percentile of degree centrality (borgatti2005centrality). This step eliminated peripheral hashtags with minimal connections to the broader substance use discourse, removing those that were weakly connected or contextually incidental rather than substantively related to substance use discussions. Next, we applied k-core decomposition ($k=10$) to identify the densely connected core of the network (Seidman and Foster, 1978). K-core decomposition recursively removes nodes with degrees less than $k$ until all remaining nodes have at least $k$ connections, ensuring a cohesive network structure. The value of $k=10$ was selected based on exploratory analysis, which indicated that this threshold preserved network connectivity while excluding sparsely connected hashtags. The resulting refined network of hashtags contained 2,333 nodes and 46,990 edges, representing the most interconnected hashtags in substance use-related content.
@@
Table 1. Categories for semantically classifying hashtags from TikTok content, their definitions, and example hashtags.
3.2. Substance Use Hashtags Network Analysis (RQ1)
3.2.1. Community Classification:
@@
We first applied established community detection algorithms, including Clauset-Newman-Moore (Clauset et al., 2004) and Louvain method (Blondel et al., 2008), but these approaches did not result in clusters that meaningfully captured substance-related discussions, as the resulting communities were driven primarily by network density rather than thematic or semantic coherence. Therefore, we used a hybrid approach combining manual and automated methods to identify thematic groups within the hashtags. We first manually annotated a sample of 251 hashtags drawn from prior work, manual TikTok searches and DEA fact sheets (United States Drug Enforcement Administration, 2025) and iteratively established substance-related content categories. This iterative process was informed by prior work on substance use discourse and refined through repeated examination of hashtag usage and contexts. As a result, we identified 16 categories representing topics related to substance use and consumption, identity, and community, health and emotion as well as awareness and advocacy. In addition, we created a Misc category for general hashtags, not related to substance use content. The final set of categories and their definitions are shown in Table 1. These categories were then used to prompt GPT-4o (OpenAI, 2024) to classify hashtags. When evaluated against human-labeled data, our best-performing prompt achieved an overall accuracy of $89.36\%$, demonstrating strong reliability in the thematic categorization. We then used this prompt to classify the full set of 2,333 hashtags. The resulting labeled hashtags were mapped back onto the network, creating themed communities that reflect different aspects of substance use discourse on TikTok.
3.2.2. Network Analysis: